OpenStudy (anonymous):

Which of the following is a factor of 8 – 125x3? 2 + 5x 4 + 20x + 25x2 4 + 10x + 25x2 None of the above

OpenStudy (anonymous):

Use this: \[a^3 - b^3 = (a-b)(a^2 + b^2 + ab)\]

OpenStudy (anonymous):

You can write the above expression as: \[2^3 - (5x)^3\] Now use the above formula..

Parth (parthkohli):

First - express it as \(2^3 - (5x)^3\) \(\Rightarrow a^3 - b^3 = (a - b)(a^2 + ab + b^2) \Longrightarrow 2^3 - (5x)^3 = (2 - 5x)(4 + 10x + 25x^2)\)

OpenStudy (anonymous):

@ waterineyes is it 4 + 10x + 25x2

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

But the entire is: \[(2−5x)(4+10x+25x^2)\]